hi i really need your help.I really dont know how to do this can you help me.
it says to evaluate:
52 ÷ (-22 + 32) + 24 · (1/4) =
 
122 - 42 ÷ (-1/2) + 2 · (-3)2 =

Answer by checkley71(8403)

52/(-22+32)+24*(1/4)=52/10+6=26/5+6=5.2+6=11.2 ANSWER.
-------------------------------------------
122-42/(-1/5)+2*(-3)2=122-42/.2+2*-6=122-210-12=100 ANSWER.